OpenStudy (anonymous):

Please, tell me what to do without giving me the answer. Perform the indicated operations and leave your answer in simplest form. Assume all variables are unrestricted except by natural domains. 1) l 3 - 4i l 2) cube root of (1/9 (x-5)^2 + 5)

OpenStudy (anonymous):

#2 looks like this \[\sqrt[3]{\frac{ 1 }{ 9 }(x -5)^{2}+5}\]

OpenStudy (anonymous):

The first one is finding the abs value or magnitude of the imaginary number The second one, you're going to use DeMoivre's formula to solve the equation Ready?

OpenStudy (anonymous):

So, for the first one, I have to find the distance of the real and imaginary part of it? Would the real part be 3 and the imaginary part 4i? And for the second one, what is DeMoivre's formula? We never went over that in class so its new to me.

OpenStudy (anonymous):

Sorry, I used the wrong method for the second one, we'll get back to that.

OpenStudy (anonymous):

For the first one, to find the magnitude of a complex number (a number made up of an imaginary and a real part real: has no i on it imaginary: has an i on it where i=\[\Large \sqrt{-1}\])

OpenStudy (anonymous):

The formula is this where a and b are real number and bi is the imaginary part however\[ \Large \left| a+bi \right|=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

So you got to plug in the numbers for a and b and solve

OpenStudy (anonymous):

oh, so i would have a=3 and b=4 which makes that \[\sqrt{(3)^{2}+(4)^{2}}\] and my answer would be 25. Is that correct?

OpenStudy (anonymous):

sqrt it

OpenStudy (anonymous):

Oh, sorry. So my final answer is 5. Thank you! However, for the second one, will I multiply everything first?

OpenStudy (anonymous):

I'm trying to find out the method for doing this, I would suppose that you don't multiply it out

OpenStudy (anonymous):

I'm just gonna go with it in simplified form

OpenStudy (anonymous):

First you gotta expand the (x-5)^2

OpenStudy (anonymous):

like (x-5)(x-5) or (x^2 - 10x +25) ?

OpenStudy (anonymous):

Correct. Now I would say to multiply the whole expression you multiplied through by 1/9 but it'll be eaiser to just put everything over 9 just\[\Large \frac{ x^2-10x+25 }{ 9 }+5\] like this

OpenStudy (anonymous):

And then you would convert that 5 at the end into a number of 9 so that we can add it to the whole expression. What number would that be?

OpenStudy (anonymous):

45/9? and that makes it \[\frac{ x ^{2}-10x+25+45 }{ 9 }\] do i simplify it afterwards and then cube root it? Or am I missing a step?

OpenStudy (anonymous):

Yeah and then combine to get 70 as the constant

OpenStudy (anonymous):

And all of that will have to get 1/9 factored out (I only multiplied it to combine it with the 5 and then take it out) and you end up with \[\Large \sqrt[3]{\frac{1}{9} (x ^{2}-10x+70)}\]

OpenStudy (anonymous):

You gotta separate the 1/9 ftom the rest of the expression \[\sqrt[3]{a \times b}=\sqrt[3]{a}\times \sqrt[3]{b}\]

OpenStudy (anonymous):

\[\Large \sqrt[3]{\frac{1}{9}}\times \sqrt[3]{(x ^{2}-10x+70)}\]

OpenStudy (anonymous):

The Right Hand Side can be simplified to \[\Large (x ^{2}-10x+70)^\frac{1}{3}\]

OpenStudy (anonymous):

do i do the same thing for the left hand side?

OpenStudy (anonymous):

The cbrt(1/9) can be rationalized to \[\Large \sqrt[3]{\frac{ 1 }{ 9 }}=\frac{ 1}{ \sqrt[3]{9} } \times \frac{ \sqrt[3]{9} }{ \sqrt[3]{9} } \times \frac{ \sqrt[3]{9} }{ \sqrt[3]{9}}\]

OpenStudy (anonymous):

\[\Large \frac{ \sqrt[3]{81} }{ 9 }\]

OpenStudy (anonymous):

\[\Large \frac{ \sqrt[3]{81} }{ 9 } \times (x ^{2}-10x+70)^\frac{1}{3}\] would be your final answer, I don't know if they want it simplified more but that's it, I gotta go, good luck!!1

OpenStudy (anonymous):

thank you! :D